# Functional Critic Modeling for Provably Convergent Off-Policy Actor-Critic

### v2

@@ Abstract
Off-policy reinforcement learning (RL) with function approximation offers an effective way to improve sample efficiency by reusing past experience. Within this setting, the actor–critic (AC) framework has achieved strong empirical success. However, both the critic and actor learning is challenging for the off-policy AC methods: first of all, in addition to the classic “deadly triad” instability of off-policy evaluation, it also suffers from a “moving target” problem, where the policy being evaluated changes continually; secondly, actor learning becomes less efficient due to the difficulty of estimating the exact off-policy policy gradient. The first challenge essentially reduces the problem to repeatedly performing off-policy evaluation for changing policies. For the second challenge, the off-policy policy gradient theorem requires a complex and often impractical algorithm to estimate an additional emphasis critic, which is typically neglected in practice, thereby reducing to the on-policy policy gradient as an approximation. In this work, we introduce a novel concept of functional critic modeling, which leads to a new AC framework that addresses both challenges for actor-critic learning under the deadly triad setting. We provide a theoretical analysis in the linear function setting, establishing the provable convergence of our framework, which, to the best of our knowledge, is the first convergent off-policy target-based AC algorithm. From a practical perspective, we further propose a carefully designed neural network architecture for the functional critic modeling and demonstrate its effectiveness through preliminary experiments on widely-used RL tasks from the DeepMind Control Benchmark.
11footnotetext: These authors contributed equally to this work.11footnotetext: Horizon Robotics. {qinxun.bai, wei.xu}@horizon.auto22footnotetext: Stern School of Business, New York University. {yh6061, zzhou}@stern.nyu.edu
@@ 1 Introduction
Challenges.
@@
Despite the empirical success of off-policy AC, this framework still faces several challenges—both in practice and in theory. One widely-studied difficulty lies in the critic update step, which conducts off-policy evaluation for the target policy. In contrast to the on-policy setting, applying standard evaluation techniques such as temporal-difference (TD) learning with function approximation under the off-policy setting can lead to instability—a phenomenon known as the “deadly triad” problem (Baird et al., 1995; Tsitsiklis and Van Roy, 1996; Kolter, 2011; Sutton and Barto, 2018). Although there are convergent off-policy evaluation methods under the deadly triad, such as gradient TD methods (Sutton et al., 2009; Qian and Zhang, 2025; Yu, 2017; Maei et al., 2009), emphatic TD methods (Sutton et al., 2016; Yu, 2015), and target critic-based methods (Mnih et al., 2015; Lee and He, 2019; Zhang et al., 2021; Chen et al., 2023), provable convergence within the off-policy AC framework remains largely unresolved, due to two additional challenges in critic learning and actor learning respectively. First, under the AC framework, the evaluated policy is constantly changing, as the actor is updated continuously—leading to a moving target problem for the off-policy evaluation/critic learning. In fact, even for on-policy AC, existing convergence guarantee depends on a two-timescale framework (Konda, 2002), where the policy has to be updated much more slowly than the updates driving the critic’s convergence. Second, the off-policy policy gradient (Imani et al., 2018) is more complex than its on-policy counterpart, as it involves not only value estimation but also correction terms, known as the emphatic term, to account for distribution mismatch. These corrections introduce additional sample-based estimation steps to critic learning, which is vulnerable to the same instability issues as the policy-evaluation step (Zhang et al., 2020).
State-of-the-art.
@@
To address the instability issue and additional challenges of off-policy actor–critic (AC), Zhang et al. (2020)—to our knowledge the only prior work with a convergence guarantee—combines two types of provably stable gradient-based evaluation methods, known as gradient TD (GTD) (Sutton et al., 2009) and Emphatic TD (ETD)(Sutton et al., 2016) within a two-timescale framework. While this ensures convergence, it substantially slows policy improvement, since most samples are consumed by critic updates rather than policy updates. Moreover, the GTD and ETD-based algorithm is complicated, requiring estimation of several additional critic-like quantities that introduce extra computational and numerical burdens as well as provable instability concerns under practical settings (Manek and Kolter, 2022). In practice, most successful off-policy RL algorithms (Fujimoto et al., 2018; Haarnoja et al., 2018; Lillicrap et al., 2015; Ball et al., 2023) use target critics to stabilize TD learning, as this approach has been empirically shown to be easier to tune than gradient TD methods. However, because of the two additional challenges noted above in off-policy AC, none of the target-based methods is provably convergent under function approximation. In particular, since off-policy policy gradients require extra sample-based estimation for the emphatic term—and the corresponding target-based estimation techniques remain unclear—target-based AC methods often fall back to on-policy gradient formulas, avoiding algorithmic complications but resulting in less efficient policy improvement and lacking the convergence guarantee of exact off-policy gradients. Moreover, to address the moving-policy issue, state-of-the-art empirical methods (Chen et al., 2021; Ball et al., 2023) typically employ a high update-to-data (UTD) ratio, which mirrors the two-timescale design but likewise reduces sample efficiency.
Our solution.
To overcome the challenges faced by existing off-policy AC methods, we introduce a new actor-critic framework with functional critics that eliminates both the need for slow-rate policy updates and the need for correction terms in exact off-policy gradients. The key idea is to model the critic as a functional that maps the policy space to policy values, for given state or state-action pair. For policy evaluation, this allows the critic to generalize to new, changing policies without needing to restart the evaluation process after every actor update. For policy improvement, exact off-policy gradients can be directly computed from the learned functional critic, without correction terms and additional estimation loops. In learning functional critics, we adopt a functional TD learning scheme that leverages the existence of the Bellman equation for each changing policy.
@@
To support our framework with theoretical insights, we analyze it in a setting under linear functional approximation—an adaptation of the standard linear value function approximation widely-used by existing theoretical work, to the functional setting (see §3.2). This simplified model enables us to rigorously demonstrate that our framework yields the first convergent off-policy target-based AC algorithm—without relying on a multi-timescale design between actor and target-critic updates (key updates driving the critic’s convergence)—thereby validating our insights of functional critic modeling and highlighting the potential of our framework to address fundamental challenges in off-policy AC.
@@
Our implementation of the functional critic leverages the expressive power of modern neural networks to approximate complex mappings between policies and value estimates (Vaswani et al., 2017; Radford et al., 2018, 2019; Dosovitskiy et al., 2020; Devlin et al., 2019). In §4, we present a minimal implementation of our proposed framework, incorporating only the theoretical insights without any bells and whistles. In two widely used continuous control tasks from the Deepmind Control Benchmark(Tassa et al., 2018), we report preliminary empirical results that compare favorably to those of a representative state-of-the-art off-policy AC method. This highlights the potential of our framework, particularly given that our implementation omits several widely adopted heuristics considered essential for existing off-policy AC methods.
@@
This work makes three key contributions,
1. The novel concept of functional critic modeling, which leads to a new AC framework that addresses the two major challenges for actor-critic learning under the deadly triad setting.
2. The first provably convergent off-policy target-based AC algorithm under function approximation.
3. A minimal implementation of the proposed AC framework using modern neural networks that achieves encouraging preliminary results.
2 Background
@@
From a practical perspective, directly ignoring the second term in (2.1) (i.e. using on-policy gradient approximation) has led to some empirical success (Degris et al., 2012; Silver et al., 2014; Lillicrap et al., 2015; Fujimoto et al., 2018) as well as local policy improvement guarantees (Degris et al., 2012). Theoretically, however, to ensure the convergence of off-policy AC, an unbiased estimation of the exact formula (2.1) is needed. Existing work relies on a policy-dependent emphasis term $m_{\theta}(s):=\mathbb{E}_{a_{t}\sim\pi_{\theta}(s_{t})}[\sum_{t=0}^{+\infty}d_{\mu}(s_{t})\lvert s_{0}=s]$, to evaluate the following exact equivalent formula (Imani et al., 2018) of (2.1),
@@
The policy-dependent emphasis term ${m}_{\theta}(s)$ reflects state dependent reweighting to correct for the sampling distribution. Like the value function, this emphasis term must be estimated from data, and is similarly prone to the instability issues mentioned above, which has to be carefully treated and combined with convergent off-policy evaluation techniques in order to have provably convergent estimates (Zhang et al., 2020). However, this additional estimation step introduces extra learning-rate schedules, which complicates hyperparameter tuning and prevents the method from scaling as effectively as prior empirical approaches, thereby creating a mismatch between empirically successful algorithms and those with provable convergence guarantees.
Finally, we conclude this section by highlighting an additional challenge in developing provably convergent algorithms for target-based off-policy AC, which is the central focus of Section 3.2. While Zhang et al. (2021); Chen et al. (2023) demonstrate that incorporating target network–based designs can provably resolve the off-policy $Q$-evaluation problem, their policy improvement guarantees are established only value-based control rather than for actor–critic. The key difficulty is that an estimator of $Q$ values alone is insufficient for computing the off-policy policy gradient in (2.1), due to the emphasis term. This limitation restricts the analysis of target-based AC algorithms to the on-policy setting (Barakat et al., 2022). This motivates us to develop a functional-approximation–based formulation, under which the off-policy gradient can be computed solely from the estimation of the $Q$ function.
@@ 3 Methodology
In this section, we introduce our off-policy actor-critic framework. §3.1 presents a general meta algorithm based on function approximations of actor and functional critic. We then investigate the theoretical guarantees under linear functional approximation in §3.2, describing the key sub-routines in detail and discussing possible alternatives that can be integrated into the meta-algorithm. In §4, we provide a minimal neural network-based implementation and preliminary experimental results.
@@ 3.1 Meta Algorithm: Functional Actor Critic
The overall procedure of our framework is presented as a meta-algorithm composed of modular sub-routines for policy evaluation (corresponding to the functional critic update) and policy improvement (corresponding to the actor update), as summarized in Algorithm 1.
@@
1:Inputs: Initial actor parameter $\theta_{0},$ initial functional critic parameter $\xi_{0},$ number of epochs $T$, batch size $m$, actor update step-size schedule $\{\eta_{t}\}_{t=1}^{T}$.
2:for $t=1,\dots,T.$ do
@@
We emphasize a key motivation behind our functional evaluator: to decouple policy optimization from value estimation while enabling the critic to generalize across time-varying input policies. Unlike standard actor-critic methods that rely on a two-timescale schedule to chase a “moving target”—where the critic must be updated many times for each actor update, and then re-updated after each actor change—our functional critic is designed to accumulate knowledge across policies in an explicit manner. The idea behind this design is that the value functions of continually changing policies often share common structures that can be effectively captured by sufficiently expressive functionals—such as those parameterized by large-scale neural networks.
@@
The actor update sub-routine in our framework follows the policy gradient paradigm. However, a key distinction lies in how we compute the exact off-policy gradient: it can be derived directly from the learned functional critic $\hat{Q}(\pi_{\theta_{t}},s,a;\xi)$,
@@
Compared with prior approaches, as discussed in §2 around (2.1), our formulation neither drops the second term as most empirical approaches, which introduces additional errors, nor relies on emphasis estimation as in Imani et al. (2018); Zhang et al. (2020), which brings further instability challenges. The simplicity of this exact off-policy gradient formula, again, shows the power of our proposed functional critic modeling.
3.2 Theoretical Benefits: A Case Study under Linear Functional Approximation

### v3

# Functional Critics Are Essential in Off-Policy Actor-Critic: Provable Convergence and Efficient Exploration

@@ Abstract
Off-policy reinforcement learning (RL) with function approximation offers an effective way to improve sample efficiency by reusing past experience. Within this setting, the actor–critic (AC) framework has achieved strong empirical success. However, this framework suffers from the ”moving target” problem, where the policy being evaluated changes continually (Dabney et al., 2021). Functional critics, also known as policy-conditioned value functions (Harb et al., 2020; Tang et al., 2022; Faccio et al., 2020), have been proposed to address this issue by explicitly including a representation of the policy as input. While the concept of generalizing value functions across the space of policies is appealing, previous efforts have struggled to remain competitive against state-of-the-art actor-critic algorithms that do not utilize functional critics (Griesbach, 2023). In this work, we revisit the concept of functional critics within the off-policy actor-critic framework. In particular, we identify two critical aspects that render functional critics a necessity rather than a luxury. First, in off-policy AC methods, critic learning contends with both the classic “deadly triad” instability and the “moving target” issue, while actor learning faces the challenge of estimating the exact off-policy policy gradient. The complex interplay between these factors renders theoretical convergence, even under linear function approximations, extremely difficult for practical off-policy AC algorithms. We demonstrate that a functional critic is essential for addressing this challenge and establish, to the best of our knowledge, the first convergence proof for an off-policy target-based AC algorithm under linear function approximation. Second, we identify a crucial link between functional critic modeling and the challenge of efficient exploration in RL. Specifically, we point out that approximating posterior sampling for exploration in a model-free setting is infeasible without the use of functional critics. From a practical perspective, we propose a tailored neural network architecture for the functional critic, along with a minimal actor-critic algorithm that relies solely on the insights above, without the need for standard implementation heuristics. In preliminary experiments on the DeepMind Control Suite, this minimal implementation achieves performance competitive with state-of-the-art actor-critic methods.
11footnotetext: These authors contributed equally to this work.22footnotetext: Work done at Horizon Robotics.11footnotetext: Isara Laboratories. jerry@isara.io22footnotetext: Stern School of Business, New York University. {yh6061, zzhou}@stern.nyu.edu33footnotetext: Horizon Robotics. wei.xu@horizon.auto
@@ 1 Introduction
Despite the empirical success of off-policy AC, this framework still faces several challenges—both in practice and in theory. Two of the arguably most critical challenges are training stability and efficient exploration. The latter concerns the issue of acquiring the most useful experience from the environment, while the former concerns how to effectively train the policy given such experience.
On the other hand, the concept of a Functional Critic (or Policy-Conditioned Value Function) has been proposed (Harb et al., 2020; Tang et al., 2022; Faccio et al., 2020) to address the difficulty caused by the “moving target” problem in AC—where the evaluated policy changes continuously as the actor is updated. The core idea of functional critic is elegant: rather than relearning a transient value function for every new policy, the critic learns a global functional $Q(s,a,\pi)$ that takes the policy representation as an explicit input. Despite the intuitive appeal of functional critics, it remains unclear whether they can effectively address the critical challenges of training stability and efficient exploration. Prior works remain mostly heuristic in arguing the effectiveness of functional critics, and a rigorous understanding of the relation between functional critics and the convergence properties of off-policy AC is missing from the literature. Note that convergence analysis has been explicitly listed under future work in Faccio et al. (2020) and Tang et al. (2022). Furthermore, the relation between functional critics and efficient exploration in off-policy AC is completely underinvestigated in the literature. Consequently, it is not surprising that previous efforts in functional critics have failed to achieve competitive empirical performance against state-of-the-art AC algorithms (Griesbach, 2023).
In this work, we uncover the deeper significance of functional critic modeling. We demonstrate that while functional critics were originally motivated by the “moving target” problem, they are effectively the missing piece required to solve the broader challenges of stability and exploration in off-policy AC.
@@
Regarding training stability, the central difficulty stems from the interplay between the “deadly triad” and the “moving target” problem. In contrast to the on-policy setting, applying standard evaluation techniques such as temporal-difference (TD) learning with function approximation under the off-policy setting can lead to instability—a phenomenon known as the “deadly triad” (Baird and others, 1995; Tsitsiklis and Van Roy, 1996; Kolter, 2011; Sutton and Barto, 2018). Although there are convergent off-policy evaluation methods designed to withstand the deadly triad, such as gradient TD methods (Sutton et al., 2009; Qian and Zhang, 2025; Yu, 2017; Maei et al., 2009), emphatic TD methods (Sutton et al., 2016; Yu, 2015), and target critic-based methods (Mnih et al., 2015; Lee and He, 2019; Zhang et al., 2021; Chen et al., 2023), provable convergence within the off-policy AC framework remains largely unresolved. This is due to two additional challenges in critic learning and actor learning respectively. For critic learning, the “moving target” problem exists even in on-policy AC, and existing convergence guarantees depend on a two-timescale framework (Konda, 2002), where the policy must be updated much more slowly than the critic to ensure convergence. For actor learning, the off-policy policy gradient (Imani et al., 2018) is more complex than its on-policy counterpart, as it involves not only value estimation but also correction terms (e.g., emphatic weights) to account for distribution mismatch. These corrections introduce additional sample-based estimation steps to critic learning, rendering it vulnerable to the same instability issues as the policy-evaluation step (Zhang et al., 2020).
@@
Zhang et al. (2020)—to our knowledge the only prior work with a convergence guarantee for off-policy AC—combines two types of provably stable gradient-based evaluation methods, Gradient TD (GTD) (Sutton et al., 2009) and Emphatic TD (ETD) (Sutton et al., 2016), within a two-timescale framework. However, this approach is complex, requiring the estimation of several additional critic-like quantities that introduce extra computational and numerical burdens, as well as potential instability concerns in practical settings (Manek and Kolter, 2022). In contrast, most successful practical off-policy RL algorithms (Fujimoto et al., 2018; Haarnoja et al., 2018; Lillicrap et al., 2015; Ball et al., 2023) utilize target critics to stabilize TD learning, as this approach has been empirically shown to be easier to tune than gradient-based methods. However, due to the aforementioned challenges in off-policy AC, none of these target-based methods are provably convergent under function approximation.
@@
In this work, we bridge this gap by revisiting the functional critic framework. In particular, we analyze the actor-critic framework with functional critics in a setting under linear functional approximation—an adaptation of the standard linear value function approximation widely-used by existing theoretical work to the functional setting (see §3.2). This simplified model enables us to rigorously demonstrate that the functional critic-based framework yields the first convergent off-policy target-based AC algorithm—without relying on a multi-timescale design between actor and target-critic updates (key updates driving the critic’s convergence)—thereby validating the theoretical motivation of functional critics and highlighting the necessity of functional critics in ensuring convergent off-policy AC.
Regarding efficient exploration, Posterior Sampling Reinforcement Learning (PSRL) stands out as a strong model-based approach, originally inspired by the Bayesian optimal Thompson Sampling algorithm for multi-arm bandits. However, due to its computational intractability in high-dimensional settings, various model-free approximations of PSRL have been proposed in the literature (Osband et al., 2016, 2018, 2019). We demonstrate the fundamental limitations of these previous approximations and illustrate that functional critics are the right tool to overcome these limitations, enabling a true posterior sampling style of exploration for actor-critic methods.
@@
To validate these theoretical insights, we propose a concrete implementation of the functional critic that leverages the expressive power of modern neural networks to approximate complex mappings between policies and value estimates (Vaswani et al., 2017; Radford et al., 2018, 2019; Dosovitskiy et al., 2020; Devlin et al., 2019). In §4, we present a minimalist implementation of this framework, incorporating only the core principles regarding the necessity of functional critics for stability and exploration. We report preliminary empirical results on widely used continuous control tasks from the Deepmind Control Suite(Tassa et al., 2018). Our method compares favorably to representative state-of-the-art off-policy AC algorithms, highlighting the potential of our framework—particularly given that our implementation omits several widely adopted heuristics often considered essential for performance.
@@
This work makes three key contributions:
  1. 1.
    Stability & Convergence: We provide, to the best of our knowledge, the first convergence proof for an off-policy target-based AC algorithm under linear function approximation. We show that the functional critic is the key element that enables this result by resolving the “deadly triad” and “moving target” instabilities.
  2. 2.
    Efficient Exploration: We reveal new connections between functional critics and exploration in model-free RL by demonstrating that approximating posterior sampling-based exploration in a model-free setting is infeasible without the functional critic formalism.
  3. 3.
    A Minimalist Algorithm: Translating our theory into practice, we propose a tailored neural network architecture and a minimalist algorithm. In experiments on the DeepMind Control Suite, this simple framework achieves performance competitive with state-of-the-art methods.
2 Preliminaries
@@
There exist two formulas for exact off-policy policy gradient. This first relies on a policy-dependent emphasis term $m_{\theta}(s):=\mathbb{E}_{a_{t}\sim\pi_{\theta}(s_{t})}[\sum_{t=0}^{+\infty}d_{\mu}(s_{t})\lvert s_{0}=s]$, to evaluate the following exact equivalent formula (Imani et al., 2018) of (2.1),
@@
The policy-dependent emphasis term ${m}_{\theta}(s)$ reflects state dependent reweighting to correct for the sampling distribution. Like the value function, this emphasis term must be estimated from data, and is similarly prone to the instability issues mentioned above. This additional estimation step introduces extra learning-rate schedules, which complicates hyperparameter tuning and prevents the method from scaling as effectively as simpler approaches.
The second formula (Faccio et al., 2020) leverages the functional critic to directly estimate (2.2),
where $\hat{Q}(\pi_{\theta},s,a;\xi)$ is the functional critic parameterized by $\xi$. In this work, we revisit functional critics and therefore use (2.3) for actor update.
Posterior Sampling for Reinforcement Learning (PSRL).
PSRL is inspired by a classical Bayesian approach to the multi-arm bandit problem known as Thompson Sampling (TS). TS maintains a posterior of the expected reward for each arm and selects an arm according to the probability that it is the optimal one. This algorithm achieves optimal Bayesian regret by ensuring efficient exploration while minimizing long-term regret. PSRL extends this Bayesian idea to the more complex RL setting. It is a model-based approach that achieves strong performance both theoretically and empirically in tabular setting (Osband et al., 2013). This algorithm maintains a posterior over the underlying MDP and operates in an episodic fashion: at the start of each episode, it samples a single MDP from the posterior, computes the optimal policy for the sampled MDP, and follows this policy for the whole episode. The collected experience is then used to update the posterior over MDP.
@@ 3 Methodology
In this section, we introduce our off-policy actor-critic framework and discuss its advantages. §3.1 presents a general meta algorithm based on function approximations of actor and functional critic. We then investigate the first critical challenge, training stability, in §3.2. We describe key sub-routines in detail and show convergence guarantees under linear functional approximation. The second critical challenge, efficient exploration, is discussed in §3.3. We demonstrate the fundamental limitations of existing model-free approximations of PSRL and illustrate that functional critics are the key missing tool to address these limitations. Following our analysis, an extension of the meta algorithm is presented to realize the potential in posterior sampling-style of exploration.
@@ 3.1 Meta Algorithm: Functional Actor Critic
The overall procedure of the functional critic framework is presented as a meta-algorithm composed of modular sub-routines for policy evaluation (corresponding to the functional critic update) and policy improvement (corresponding to the actor update), as summarized in Algorithm 1.
@@
1:Inputs: Initial actor parameter $\theta_{0},$ initial functional critic parameter $\xi_{0},$ number of epochs $T$, batch size $m$, actor update step-size schedule $\{\eta_{t}\}_{t=1}^{T}$
2:for $t=1,\dots,T$ do
@@
We emphasize a key motivation behind functional evaluator: to decouple policy optimization from value estimation while enabling the critic to generalize across time-varying input policies. The idea behind this design is that the value functions of continually changing policies often share common structures that can be effectively captured by sufficiently expressive functionals—such as those parameterized by large-scale neural networks.
@@
The actor update sub-routine in our framework follows the off-policy policy gradient proposed in Faccio et al. (2020), where the exact off-policy gradient can be derived directly from the learned functional critic $\hat{Q}(\pi_{\theta_{t}},s,a;\xi)$,
@@
3.2 Convergence Guarantee: A Case Study under Linear Functional Approximation
@@
Remark 3.
It is worth noted that functional critic is necessary for developing provably convergent algorithms for target-based off-policy AC, which is the central focus of this section. The rationale lies in two folds. Firstly, from a practical perspective, directly ignoring the second term of (2.1) (i.e. using on-policy gradient approximation) in actor learning has led to some empirical success (Degris et al., 2012; Silver et al., 2014; Lillicrap et al., 2015; Fujimoto et al., 2018) as well as local policy improvement guarantees (Degris et al., 2012). Theoretically, however, to ensure the convergence of off-policy AC, an unbiased estimation of the exact off-policy gradient (2.1) is needed, which depends on either (2.2) or (2.3). Secondly, it is unclear how to combine the empirically successful target-based TD-learning with the off-policy gradient formula (2.2) to form a convergent off-policy AC algorithm, due to the challenges of estimating the extra emphasis term as explained above. As a result, existing convergent target-based methods are either for value-based control  (Zhang et al., 2021; Chen et al., 2023) or for on-policy AC (Barakat et al., 2022) only.
3.3 Efficient Exploration: Model-Free Approximation of PSRL
As reviewed in §2, in multi-arm bandits, Thompson Sampling (TS) maintains a posterior of reward for each arm, as shown in Figure LABEL:fig:TS. Extending to model-based RL, as shown in Figure LABEL:fig:RL, the return depends on both the underlying MDP and the executed policy. PSRL extends the idea of TS to model-based RL by maintaining a posterior distribution over the underlying MDP. For each rollout, it first samples an MDP from the posterior, which slices the hyper-surface in Figure LABEL:fig:RL to get the curve in Figure LABEL:fig:mdp_opt_pi, and extracts the optimal policy from the peak of the curve for rollout.
Following the above intuition, a model-free approximation of PSRL should maintain a distribution over the set of “green curves” of Figure LABEL:fig:mdp_opt_pi. Existing approaches (Osband et al., 2016, 2018, 2019), however, end up keeping an ensemble of standard critics, which does not capture any sense of “distribution over the curves”. To see this, let us discuss respectively the two possible approaches of constructing target critics for training the critic ensemble.
The first approach, as shown in Figure LABEL:fig:ensemble_q_same_tar, adopts a same target critic for all critics of the ensemble. In this way, the ensemble of critics are different estimates of the same quantity, i.e., $q^{\hat{\pi}}$ for the target policy $\hat{\pi}$ at that time. As a result, the ensemble of critics, at its best, can only model the posterior of $q^{\hat{\pi}}$ for one fixed $\hat{\pi}$ at a time. It cannot capture any distribution over the whole green curve and therefore does not enable posterior sampling-style of exploration.
The second approach, as shown in Figure LABEL:fig:ensemble_q_diff_tar, adopts a different target critic for each individual critic of the ensemble. In this way, the ensemble of critics are independent estimates of different quantities, i.e., $q^{\hat{\pi}_{i}}$ for different target policies $\hat{\pi}_{i}$. As a result, the ensemble of critics, at its best, can only model a small number of sampled points of a green curve, rather than any distribution over the whole green curves. It does not enable posterior sampling-style of exploration either.
What we actually need is an ensemble of functional critics, where each member exactly corresponds to an individual estimate of the green curve. And therefore the ensemble serves as a particle-based approximation of the posterior over green curves. Motivated by this insight, we extend the meta Algorithm 1 to employ an ensemble of functional critics and pair each of them with corresponding actors, as detailed in Algorithm 3.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 1: Ensemble of critics do not approximate the PSRL while ensemble of fuctional critics is the tool we need.
Algorithm 3 Functional Actor-Critic Algorithm with Ensemble of Actor-Critic Pairs
1:Inputs: Initial actor parameter $\big\{\theta^{(i)}_{0}\big\}^{n}_{i=1},$ initial functional critic parameter $\big\{\xi^{(i)}_{0}\big\}^{n}_{i=1},$ number of epochs $T$, batch size $m$, actor update step-size schedule $\{\eta_{t}\}_{t=1}^{T}$
2:for $t=1,\dots,T$ do
4:  Update functional critic parameter $\xi^{(i)}_{t}\leftarrowFunctional$ Policy Evaluator ($t,\xi^{(i)}_{t-1},\theta^{(i)}_{t-1},\mathcal{D}_{t}$) for each $i$
5:  Compute the Parameterized Off-Policy Gradient $G^{(i)}_{t}\leftarrow\nabla_{\theta}\big(\mathcal{J}(\pi_{\theta};\xi^{(i)}_{t})\big)\lvert_{\theta=\theta^{(i)}_{t-1}}$ for each $i$
6:  Update the actor parameter $\theta^{(i)}_{t}\leftarrow\theta^{(i)}_{t-1}-\eta_{t}G^{(i)}_{t}$ for each $i$
7:end for
8:Return $\theta_{T}$